OpenStudy (anonymous):

Prove the trig. identity: ((sinx)/(1+cosx))+((1+cosx)/(sinx))=2cscx

OpenStudy (anonymous):

best bet is to add. if it is not clear how to add this stuff up, replace sin(x) by a and cos(x) by b and add \[\frac{b}{1+a}+\frac{a}{1+b}\]

OpenStudy (anonymous):

LHS=((sinx)/(1+cosx))+((1+cosx)/(sinx)) LHS=(sin^2x+(1+cosx)(1+cosx))/((1+cosx)(sinx))

OpenStudy (anonymous):

the race is on!

OpenStudy (anonymous):

after you add, replace back. then i bet the only "trig" step will be to write \[\sin^2(x)+\cos^2(x)=1\] and you will get your answre

OpenStudy (anonymous):

LHS=(sin^x+cos^x+2sincosx+1)/(1+cosx)sinx

OpenStudy (anonymous):

lol @ joe

OpenStudy (anonymous):

LHS=(2+2sincosx)/(1+cosx)sinx

OpenStudy (anonymous):

i either win or lose, depending on whether you think i actually wrote an answer. @kbix, most of these problems just require a bunch of algebra. don't let the "sine" and "cosine" fool you or make it more difficult. they are usually cooked up so that at some point you will see \[\sin^2(x)+\cos^2(x)\] which you then replace by 1

OpenStudy (anonymous):

just combine the two fractions: (sinx / 1 + cosx) + (1 + cosx / sinx) =sin^2(x)/ sin(x) *(1 +cos(x))^2/ sin(x) *(1 +cos(x)) =sin^2(x) +(1 +cos(x))^2/ sin(x) *(1 +cos(x)) =sin^2(x) +cos^2(x) +2cos(x) +1/ sin(x) *(1 +cos(x)) =1 +2cos(x) +1/ sin(x) *(1 +cos(x)) =2 +2cos(x) / sin(x) *(1 +cos(x)) =2 *(1 +cos(x)) / sin(x) *(1 +cos(x)) =2/ sin(x) =2csc(x)

OpenStudy (anonymous):

LHS=2 *(1 +cos(x)) / sin(x) *(1 +cos(x))

OpenStudy (anonymous):

2/sinx

OpenStudy (anonymous):

i feel cheated

OpenStudy (anonymous):

\[\frac{b}{1+a}+\frac{1+a}{b}=\] \[=\frac{b^2+(1+a)(1+a)}{(1+a)b}\] \[=\frac{b^2+a^2+2a+1}{(1+a)b}\]

OpenStudy (anonymous):

\[\frac{\sin^2(x)+\cos^2(x)+2\cos(x)+1}{(1+\cos(x))\sin(x)}\] \[\frac{2+2\cos(x)}{(1+\cos(x))\sin(x)}\] that is the ONLY trig step. the rest is algebra pure and (not so) simple

OpenStudy (anonymous):

now factor out the 2, cancel the 1 + cos(x) and you get your answer.

OpenStudy (anonymous):

@lagrange thanks. i am very slow at latex tonight

OpenStudy (anonymous):

Thank-you soooo much, guys. I am pretty much awful at proofs. I greatlyyyyy appreciate it. (: